Four years ago, the sum of the ages of Bill and Jane was 45.

Six year from now, twice jane’s age will be 28 more than

Bill’s age then. How old is jane now?


1
Expert's answer
2021-09-27T16:17:51-0400

Four years ago ;

Bill (x) ="x-4"

Jane (Y) ="y-4"

Sum of their ages =x+y-8=45 (x+y=53).........(i)


Six years from now ;

"2y-x=28"...........(ii)

Combining equation (i) and (ii) by substitution ;

"x+y=53"

"2y-x=28"

"2y-53+y=28"

3y=81

y=27

Thus x=26 and y=27

Jane's age now is 27yrs